# Energy-Efficient Routing Protocol in Vehicular Opportunistic Networks: A Dynamic Cluster-based Routing Using Deep Reinforcement Learning

### v3

@@ IV Proposed method
Clustering in OppNets organizes vehicles into groups based on parameters such as distance and location to enhance information exchange and communication. Various algorithms optimize this process [A29], [A30], [A31], [A32], but selecting an effective cluster head either centrally via infrastructure or distributively among vehicles remains a key challenge [A33]. In dynamic vehicular networks, traditional clustering struggles with rapid topology changes, causing packet loss and routing inefficiencies. RL offers a solution by enabling adaptive, real-time decision-making. To address these challenges, we propose CR-DRL, an AC based clustering and routing approach designed to optimize communication efficiency in OppNets. CR-DRL approach leverages the strengths of the AC algorithm to enhance the efficiency and effectiveness of cluster head selection. Initially, we compute the Manhattan distance between each node and its neighbors using GPS data. We assume that $d$ denotes the Manhattan distance between node $v_{i}$ and node $v_{j}$ (${d}={|{x}_{i}-{x}_{j}|}+{|{y}_{i}-{y}_{j}|}$). Subsequently, through the definition of three states involving the highest encounter rates (${\phi}_{{v_{i}}_{\max}}$), remaining energy levels ($\Psi_{{v}_{i}}$), and buffer size ($\eta_{{v}_{i}}$), the selection of candidate cluster heads will be determined.
@@ V Simulation setting and evaluation metrics
This simulation study evaluates the performance of the proposed CR-DRL protocol in dynamic OppNets. Specifically, we analyze its ability to form stable clusters and maintain efficient data routing under varying network conditions. We evaluate CR-DRL’s performance in relation to mobility patterns and vehicle movement using the Opportunistic Network Environment (ONE) simulator. ONE is an open-source Java-based simulation tool for developing and testing routing protocols [A34]. With ONE, it is easy to integrate contact records, route modules, applications, and reports. For the CR-DRL, we have used the default ONE simulation map from Helsinki, Finland. The geographical space for this experiment is ${4000*3500m}^{2}$ in size, with 50, 100, 150, 200, and 250 nodes distributed at random. The GPS coordinates, defined by longitude and latitude, are based on geographical space. To reflect real-world heterogeneous mobility, we use a mixed mobility model where 50% of nodes follow a Random Waypoint Model (RWP), 30% follow a Gauss-Markov pattern with dynamic speed and direction changes, and 20% exhibit clustered movements based on the Working Day Movement Model (WDM) model. This setup ensures diverse and realistic node mobility, accounting for varying speeds, path choices, and encounter frequencies [A35]. Each message generated within the network has a TTL of 300 seconds and varies in size from 250 to 750 KB. The nodes have buffer capacities between 50 and 100 MB. To simulate both low and high traffic loads, the message generation interval per node was randomly adjusted between 5 and 40 seconds. To emulate realistic heterogeneous mobility, vehicles were categorized into two groups: 80% cars with speeds between 8–15 m/s and 20% service vehicles (e.g., trams) with speeds of 4–8 m/s. A failure model was incorporated where 5% of nodes were randomly deactivated between times 1800–2400 seconds to simulate vehicle dropouts or hardware failures. The simulation parameters are set for a duration of 1 hour, with an initial energy of 4800 Joules, scanning energy of 0.92 mW/s, transmission energy of 0.08 mW/s, receive energy of 0.08 mW/s, an energy recharging interval of 2800 seconds, common members transmit range 100M and cluster head transmit range 300M. As mentioned in the section III, if a vehicle is unable to communicate with its cluster head due to being out of range, it momentarily increases its transmission range to restore connectivity. Initial experiments indicated that a 20% expansion in transmission range is typically sufficient to maintain communication in most scenarios. However, since this adjustment occurs only when required and has a minimal impact on overall network performance, it was excluded from the primary simulations to maintain focus on key metrics such as clustering efficiency, energy consumption, and routing performance. The simulation employs the ONE simulator’s standardized mobility model derived from the Helsinki map. Figure 3 presents the convergence characteristics of the CR-DRL framework during the learning process.
@@
The definition of key metrics such as delivery ratio, throughput, delay is as follows: The delivery ratio is to measure the proportion of data packets successfully delivered to their destination out of the total sent [A36]. E2E delay measures the time a packet takes to travel from source to destination across a network [A37]. A hop count indicates the total number of intermediate devices, such as nodes, that a data packet passes through from its source to its destination. Each node along the data path counts as a hop that facilitates the transfer of data from one point to another [A38]. Throughput indicates the volume of data in bits transferred from a source to a destination per unit of time (seconds) over a communication link [A39]. Residual energy represents the remaining battery capacity of each node, calculated as the initial energy minus the cumulative consumption from operations such as scanning, transmission, and reception. Normalized performance scales key metrics including delivery ratio, delay, and energy consumption between 0 and 1 to enable fair comparison [A36]. Convergence rate quantifies the speed and stability of policy improvement in the RL algorithm, assessed through reward stabilization and loss reduction over successive training epochs.
@@ V-B First scenario
The stability of cluster heads and members in CR-DRL is evaluated using three metrics cluster head lifetime, cluster member lifetime, and cluster head change rate to ensure efficient routing. Stable clusters minimize disruptions, reduce re-clustering overhead, and enhance data transmission reliability, with longer cluster head lifetimes indicating balanced resource allocation and lower change rates reflecting steady clusters, reduced energy use, and improved packet delivery efficiency. A recent paper dealing with stability of cluster heads and members is ECRDP [A41]. As mentioned in the related work section, this method uses a machine learning approach with dynamic clustering and is a suitable candidate for comparison with the CR-DRL method. Figures, 6, 7, and 8 illustrate that CR-DRL surpasses ECRDP in common member lifetime, cluster head lifetime, and cluster head change rate, particularly in high-density scenarios where CR-DRL shows improvements of over 14%, 7%, and 21%, respectively. The proposed CR-DRL approach achieves a higher cluster head lifetime and a lower cluster head change rate primarily due to the integration of the AC algorithm and the ADT mechanism. The AC algorithm enables the dynamic selection of cluster heads based on critical metrics such as encounter history, residual energy, and buffer capacity. This targeted selection ensures that nodes with higher stability and optimal resource availability are chosen as cluster heads, leading to prolonged cluster head lifetimes. Furthermore, the ADT mechanism adjusts clustering thresholds according to network density, ensuring clusters are neither too large nor too sparse. In high-density scenarios, smaller clusters reduce overhead, while in low-density conditions, larger clusters maintain connectivity. This dynamic adaptability minimizes the frequency of re-clustering events, thereby reducing the cluster head change rate.
@@
Also, in this scenario, we conducted a comprehensive performance evaluation of the proposed CR-DRL protocol in comparison with the benchmark EEHCHR protocol [New1]. To ensure a rigorous and equitable comparison, the evaluation was conducted using several key performance indicators: A round corresponds to a complete simulation cycle in which clustering, data transmission, and energy updates are executed across all nodes. First Node Dies (FND) denotes the round at which the first node depletes its energy, serving as an indicator of the network’s stability period. Half of the Nodes Die (HND) represents the round when 50% of the deployed nodes have exhausted their energy, thus reflecting the broader network lifetime. The metric Alive Nodes quantifies the percentage of nodes that remain functional as the simulation progresses, while Residual Energy indicates the proportion of remaining total network energy relative to the initial energy budget.
@@ V-C Second scenario
In the second scenario, as noted above in Section I, the CR-DRL approach allows TCP/IP routing protocols to be compatible with OppNets. Through adaptive clustering, the CR-DRL method can provide a stable connection between nodes that this connection is suitable for the TCP/IP protocol. We compare CR-DRL with SCF-Architecture and then compare the CR-DRL method with existing machine learning routing protocols defined in sources such as Mobility Adaptive [S1], RLProph [S2], and Latency‐Aware [S3]. The evaluation parameters are average Hop count, average Delivery ratio, Throughput, E2E delay and average remaining energy.
@@ V-D Experimental results
Figure 12 illustrates the number of partitions in the network over time, demonstrating the improved performance of CR-DRL regarding network stability. Both CR-DRL method and SCF-Architecture have high partitioning at the initial point, almost with 9 and 10 partitions, respectively. However, over time, CR-DRL significantly reduces partitioning and achieves a connected network with only 1 partition at 35 seconds. On the other hand, SCF-Architecture continue to experience multiple partitions (around 4-6 partitions even after 50 seconds) and hence are causing the data transfer process get seriously hindered. CR-DRL reduces partitions in the network by 85% compared to SCF-Architecture. Hence, the CR-DRL can maintain long-term connections and prevent network breakdowns, which can be suitable for TCP/IP networks.
@@
Throughput performance, depicted in Figure 18, confirms the cumulative advantage of CR-DRL. At 250 nodes, throughput improves by 3% over Mobility Adaptive, 12% over RLProph, and 7% over Latency-Aware routing. In contrast, RLProph is hindered by computational overhead that degrades performance, Mobility Adaptive protocol incurs packet losses arising from unstable routing paths under high vehicular mobility and rapidly varying topologies [S1], while Latency-Aware approach suffers congestion as traffic intensity increases.
@@ VI Conclusion and future work
[Uncaptioned image] Meisam Sharifi Sani received his M.Sc. degree in Software Engineering from Iran in 2014. He is currently pursuing his Ph.D. degree in Electrical and Telecommunication Engineering at the University of Wollongong, Australia, under the Research Training Program (RTP) scholarship. His fields of research include Intrusion Detection Systems (IDS), Internet of Things (IoT), Internet of Vehicle (loV), Machine Learning, VDTN, and Wireless Networks.
[Uncaptioned image] Saeid Iranmanesh received the B.E. and M.S. degrees in computer science from Iran, and the Ph.D. degree in electrical and telecommunication engineering from the University of Wollongong, Wollongong, NSW, Australia, in 2015. He is currently a senior Research fellow with the School of Electrical, Computer and Telecommunication Engineering, University of Wollongong. He is a senior member of IEEE and Fellow of Engineers Australia. His research interests include wireless networks, vehicular networks, intelligent transportation systems, and smart cities.
[Uncaptioned image] Raad Raad received a Bachelor of Engineering degree (Hons.1) in Electrical Engineering from the University of Wollongong, Australia in 1997, and a Ph.D. degree entitled Neuro-Fuzzy Logic Admission Control in Cellular Mobile Networks, in 2006. Since 2004, he has been with the School of Electrical, Computer and Telecommunications Engineering, University of Wollongong and he is the current Head of School. His current research interests include wireless communications, CubeSat, the IoT, and Antenna design in addition to advanced power systems. He is part of the ARC ITTC for Future Grids and ARC hub for Connected Health Sensors, both major research initiatives.
[Uncaptioned image] Faisel Tubbal received his B.E. in Electronic Technology from Tripoli, Libya (2004), and M.S. degrees in Telecommunication Engineering (2012) and Engineering Management (2013), as well as a Ph.D. in Telecommunication Engineering (2017) from the University of Wollongong, Australia. He began his career at the Libyan Centre for Remote Sensing and Space Science and later joined the University of Wollongong, where he is currently Laboratory Manager and Work Integrated Learning Coordinator. Since 2017, he has also served as Unit Convenor at Western Sydney University. He has authored a book, three chapters, and over 65 publications. His research focuses on CubeSat, wearable, and medical antennas, metamaterials, metasurfaces, and wireless power transfer. He is a recipient of multiple teaching excellence awards, including OCTAL, and a WATTLE fellow.

### v4

@@ IV Proposed method
Clustering in OppNets organizes vehicles into groups based on parameters such as distance and location to enhance information exchange and communication. Various algorithms optimize this process [25], [4], [21], [8], but selecting an effective cluster head either centrally via infrastructure or distributively among vehicles remains a key challenge [3]. In dynamic vehicular networks, traditional clustering struggles with rapid topology changes, causing packet loss and routing inefficiencies. RL offers a solution by enabling adaptive, real-time decision-making. To address these challenges, we propose CR-DRL, an AC based clustering and routing approach designed to optimize communication efficiency in OppNets. CR-DRL approach leverages the strengths of the AC algorithm to enhance the efficiency and effectiveness of cluster head selection. Initially, we compute the Manhattan distance between each node and its neighbors using GPS data. We assume that $d$ denotes the Manhattan distance between node $v_{i}$ and node $v_{j}$ (${d}={|{x}_{i}-{x}_{j}|}+{|{y}_{i}-{y}_{j}|}$). Subsequently, through the definition of three states involving the highest encounter rates (${\phi}_{{v_{i}}_{\max}}$), remaining energy levels ($\Psi_{{v}_{i}}$), and buffer size ($\eta_{{v}_{i}}$), the selection of candidate cluster heads will be determined.
@@ V Simulation setting and evaluation metrics
This simulation study evaluates the performance of the proposed CR-DRL protocol in dynamic OppNets. Specifically, we analyze its ability to form stable clusters and maintain efficient data routing under varying network conditions. We evaluate CR-DRL’s performance in relation to mobility patterns and vehicle movement using the Opportunistic Network Environment (ONE) simulator. ONE is an open-source Java-based simulation tool for developing and testing routing protocols [9]. With ONE, it is easy to integrate contact records, route modules, applications, and reports. For the CR-DRL, we have used the default ONE simulation map from Helsinki, Finland. The geographical space for this experiment is ${4000*3500m}^{2}$ in size, with 50, 100, 150, 200, and 250 nodes distributed at random. The GPS coordinates, defined by longitude and latitude, are based on geographical space. To reflect real-world heterogeneous mobility, we use a mixed mobility model where 50% of nodes follow a Random Waypoint Model (RWP), 30% follow a Gauss-Markov pattern with dynamic speed and direction changes, and 20% exhibit clustered movements based on the Working Day Movement Model (WDM) model. This setup ensures diverse and realistic node mobility, accounting for varying speeds, path choices, and encounter frequencies [7]. Each message generated within the network has a TTL of 300 seconds and varies in size from 250 to 750 KB. The nodes have buffer capacities between 50 and 100 MB. To simulate both low and high traffic loads, the message generation interval per node was randomly adjusted between 5 and 40 seconds. To emulate realistic heterogeneous mobility, vehicles were categorized into two groups: 80% cars with speeds between 8–15 m/s and 20% service vehicles (e.g., trams) with speeds of 4–8 m/s. A failure model was incorporated where 5% of nodes were randomly deactivated between times 1800–2400 seconds to simulate vehicle dropouts or hardware failures. The simulation parameters are set for a duration of 1 hour, with an initial energy of 4800 Joules, scanning energy of 0.92 mW/s, transmission energy of 0.08 mW/s, receive energy of 0.08 mW/s, an energy recharging interval of 2800 seconds, common members transmit range 100M and cluster head transmit range 300M. As mentioned in the section III, if a vehicle is unable to communicate with its cluster head due to being out of range, it momentarily increases its transmission range to restore connectivity. Initial experiments indicated that a 20% expansion in transmission range is typically sufficient to maintain communication in most scenarios. However, since this adjustment occurs only when required and has a minimal impact on overall network performance, it was excluded from the primary simulations to maintain focus on key metrics such as clustering efficiency, energy consumption, and routing performance. The simulation employs the ONE simulator’s standardized mobility model derived from the Helsinki map. Figure 3 presents the convergence characteristics of the CR-DRL framework during the learning process.
@@
The definition of key metrics such as delivery ratio, throughput, delay is as follows: The delivery ratio is to measure the proportion of data packets successfully delivered to their destination out of the total sent [27]. E2E delay measures the time a packet takes to travel from source to destination across a network [40]. A hop count indicates the total number of intermediate devices, such as nodes, that a data packet passes through from its source to its destination. Each node along the data path counts as a hop that facilitates the transfer of data from one point to another [16]. Throughput indicates the volume of data in bits transferred from a source to a destination per unit of time (seconds) over a communication link [17]. Residual energy represents the remaining battery capacity of each node, calculated as the initial energy minus the cumulative consumption from operations such as scanning, transmission, and reception. Normalized performance scales key metrics including delivery ratio, delay, and energy consumption between 0 and 1 to enable fair comparison [27]. Convergence rate quantifies the speed and stability of policy improvement in the RL algorithm, assessed through reward stabilization and loss reduction over successive training epochs.
@@ V-B First scenario
The stability of cluster heads and members in CR-DRL is evaluated using three metrics cluster head lifetime, cluster member lifetime, and cluster head change rate to ensure efficient routing. Stable clusters minimize disruptions, reduce re-clustering overhead, and enhance data transmission reliability, with longer cluster head lifetimes indicating balanced resource allocation and lower change rates reflecting steady clusters, reduced energy use, and improved packet delivery efficiency. A recent paper dealing with stability of cluster heads and members is ECRDP [12]. As mentioned in the related work section, this method uses a machine learning approach with dynamic clustering and is a suitable candidate for comparison with the CR-DRL method. Figures 6, 7, and 8 illustrate that CR-DRL surpasses ECRDP in common members lifetime, cluster head lifetime, and cluster head change rate, particularly in high-density scenarios where CR-DRL shows improvements of over 14%, 7%, and 21%, respectively. The proposed CR-DRL approach achieves a higher cluster head lifetime and a lower cluster head change rate primarily due to the integration of the AC algorithm and the ADT mechanism. The AC algorithm enables the dynamic selection of cluster heads based on critical metrics such as encounter history, residual energy, and buffer capacity. This targeted selection ensures that nodes with higher stability and optimal resource availability are chosen as cluster heads, leading to prolonged cluster head lifetimes. Furthermore, the ADT mechanism adjusts clustering thresholds according to network density, ensuring clusters are neither too large nor too sparse. In high-density scenarios, smaller clusters reduce overhead, while in low-density conditions, larger clusters maintain connectivity. This dynamic adaptability minimizes the frequency of re-clustering events, thereby reducing the cluster head change rate.
@@
Also, in this scenario, we conducted a comprehensive performance evaluation of the proposed CR-DRL protocol in comparison with the benchmark EEHCHR protocol [23]. To ensure a rigorous and equitable comparison, the evaluation was conducted using several key performance indicators: A round corresponds to a complete simulation cycle in which clustering, data transmission, and energy updates are executed across all nodes. First Node Dies (FND) denotes the round at which the first node depletes its energy, serving as an indicator of the network’s stability period. Half of the Nodes Die (HND) represents the round when 50% of the deployed nodes have exhausted their energy, thus reflecting the broader network lifetime. The metric Alive Nodes quantifies the percentage of nodes that remain functional as the simulation progresses, while Residual Energy indicates the proportion of remaining total network energy relative to the initial energy budget.
@@ V-C Second scenario
In the second scenario, as noted above in Section I, the CR-DRL approach allows TCP/IP routing protocols to be compatible with OppNets. Through adaptive clustering, the CR-DRL method can provide a stable connection between nodes that this connection is suitable for the TCP/IP protocol. We compare CR-DRL with SCF-Architecture and then compare the CR-DRL method with existing machine learning routing protocols defined in sources such as Mobility Adaptive [37], RLProph [34], and Latency‐Aware [33]. The evaluation parameters are average Hop count, average Delivery ratio, Throughput, E2E delay and average remaining energy.
@@ V-D Experimental results
Figure 12 illustrates the number of partitions in the network over time, demonstrating the improved performance of CR-DRL regarding network stability. Both CR-DRL method and SCF-Architecture have high partitioning at the initial point, almost with 9 and 10 partitions, respectively. However, over time, CR-DRL significantly reduces partitioning and achieves a connected network with only 1 partition at 35 seconds. On the other hand, SCF-Architecture continue to experience multiple partitions (around 3-6 partitions even after 50 seconds) and hence are causing the data transfer process get seriously hindered. CR-DRL reduces partitions in the network by 85% compared to SCF-Architecture. Hence, the CR-DRL can maintain long-term connections and prevent network breakdowns, which can be suitable for TCP/IP networks.
@@
Throughput performance, depicted in Figure 18, confirms the cumulative advantage of CR-DRL. At 250 nodes, throughput improves by 3% over Mobility Adaptive, 12% over RLProph, and 7% over Latency-Aware routing. In contrast, RLProph is hindered by computational overhead that degrades performance, Mobility Adaptive protocol incurs packet losses arising from unstable routing paths under high vehicular mobility and rapidly varying topologies [37], while Latency-Aware approach suffers congestion as traffic intensity increases.